The demand and supply equations for a certain item are given by
D = 5p + 40
S= -p^2 + 30p -8
Find the equilibrium Price.
The answer I got for this one was $1.29. Is this correct? The answer options to this question are:
A)$1.29
B)$1.43
C)$1.57
D)$1.71
Thanks for all your help with this.

Answer by stanbon(75887)

D = 5p + 40
S= -p^2 + 30p -8
Find the equilibrium Price.
-------------
Equilibrium means demand equals supply.
EQUATION:
5p+40 = -p^2+30p-8
p^2-25p+48=0
p=[25+-sqrt(625-4*48]/2
p=[25+-20.8]/2
p=$2.10 or $22.90
================
Cheers,
Stan H.